OpenStudy (linna):

Why \[\LARGE 1 + 3{x^2} \ne 0\quad \quad \quad \forall x \in \mathbb{R} \] The answers I've came up with... 1. because there's no real number that when we square it would give negative... 2. Parabola doesn't touch the origin [Points (0,0) ] 3. because solving it we get: \[\LARGE 1 + 3{x^2} = 0\] \[\LARGE 3{x^2} = -1\] \[\LARGE {x^2} = -\frac13 \] \[\LARGE x = \pm \sqrt{-\frac13 } \] \[\LARGE x = \pm i\sqrt{\frac13 } \] \[\LARGE x \neq \pm i\sqrt{\frac13 } \quad \quad \forall x \in \mathbb{R}\] But none of them were true. and this still's my homework... why: \[\LARGE 1 + 3{x^2} \ne 0\quad \quad \quad \forall x \in \mathbb{R} \]

OpenStudy (anonymous):

Its because no matter what real value of x you plug into 1+3x^2, the smallest value you can get is 1. This is because:\[\forall x\in \mathbb{R}, x^2\ge 0\Longrightarrow 3x^2\ge 0 \Longrightarrow 1+3x^2\ge 1\Longrightarrow 1+3x^2>0\]If this doesnt convice you, then just look at the graph of 1+3x^2, and note that it doesnt cross the x axis, so there is no value of x that will make y = 0.

OpenStudy (linna):

But sir, that's what I was trying to tell my professors but he said no... another proof ! .. anyway thank you :)